# Examining Kempe equivalence via commutative algebra

### v1

@@ Abstract.
Kempe equivalence is a classical and important notion on vertex coloring in graph theory. In the present paper, we introduce several ideals associated with graphs and provide a method to determine whether two $k$-colorings are Kempe equivalent via commutative algebra. Moreover, we give a way to compute all $k$-colorings of a graph up to Kempe equivalence by virtue of the algebraic technique on Gröbner bases. As a consequence, the number of $k$-Kempe classes can be computed by using Hilbert functions. Finally, we introduce several algebraic algorithms related to Kempe equivalence.
@@ 1. Introduction
A $k$-coloring $f$ of a graph $G$ on the vertex set $[d]:=\{1,2,\ldots,d\}$ is a map from $[d]$ to $[k]$ such that $f(i)\neq f(j)$ for all $\{i,j\}\in E(G)$. The smallest integer $\chi(G)$ such that $G$ has a $\chi(G)$-coloring is called the chromatic number of $G$. Given a $k$-coloring $f$ of $G$, and integers $1\leq i<j\leq k$, let $H$ be a connected component of the induced subgraph of $G$ on the vertex set $f^{-1}(i)\cup f^{-1}(j)$. Then we can obtain a new $k$-coloring $g$ of $G$ by setting
$g(x)=\begin{cases}f(x)&x\notin H,\\ i&x\in H\mbox{ and }f(x)=j,\\ j&x\in H\mbox{ and }f(x)=i.\\ \end{cases}$
We say that $g$ is obtained from $f$ by a Kempe switching. Two $k$-colorings $f$ and $g$ of $G$ are called Kempe equivalent, denoted by $f\sim_{k}g$, if there exists a sequence $f_{0},f_{1},\ldots,f_{s}$ of $k$-colorings of $G$ such that $f_{0}=f$, $f_{s}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching. Let denote $\mathcal{C}_{k}(G)$ the set of all $k$-colorings of $G$. Then $\sim_{k}$ is an equivalence relation on $\mathcal{C}_{k}(G)$. The equivalence classes of $\mathcal{C}_{k}(G)$ by $\sim_{k}$ are called the $k$-Kempe classes. We denote ${\rm kc}(G,k)$ the quotient set $\mathcal{C}_{k}(G)/\sim_{k}$ and denote ${\rm Kc}(G,k)$ the number of $k$-Kempe classes of $G$, namely ${\rm Kc}(G,k)=|{\rm kc}(G,k)|$. See [5] for an overview of Kempe equivalence.
@@
Let $G$ be a graph on $[d]$ and $<$ a monomial order on $R[G]$, and let $\{{\mathbf{x}}_{f_{1}},\dots,{\mathbf{x}}_{f_{s}}\}$ be the set of all standard monomials of degree $k$ with respect to the initial ideal ${\rm in}_{<}(K_{G})$. Then
@@
The present paper is organized as follows: In Section 2, we will recall the definition of stable set ideals and explain a relationship between stable set ideals and Kempe equivalence. Section 3 will give a brief introduction on Gröbner bases. In Section 4, we will define $2$-coloring ideals and see their algebraic properties. In Section 5, a proof of Theorem 1.1 will be given. In Section 6, we will define Kempe ideals and prove Theorem 1.2. Finally, in Section 7, we will introduce several algorithms related to Kempe equivalence.
@@ 2. Stable set ideals
We regard ${\mathbb{K}}[G]$ as a homogeneous algebra by setting each $\deg({\mathbf{t}}^{\rho(S_{i})}s)=1$. Note that ${\mathbb{K}}[G]$ is a toric ring. Let $R[G]={\mathbb{K}}[x_{S_{1}},\ldots,x_{S_{n}}]$ denote the polynomial ring in $n$ variables over ${\mathbb{K}}$ with each $\deg(x_{S_{i}})=1$. The stable set ideal of $G$ is the kernel of the surjective homomorphism $\pi:R[G]\to{\mathbb{K}}[G]$ defined by $\pi(x_{S_{i}})={\mathbf{t}}^{\rho(S_{i})}s$ for $1\leq i\leq n$. Note that $I_{G}$ is a toric ideal, and hence a prime ideal generated by homogeneous binomials. The toric ring ${\mathbb{K}}[G]$ is called quadratic if $I_{G}$ is generated by quadratic binomials. We say that “$I_{G}$ is generated by quadratic binomials” even if $I_{G}=\{0\}$ (or equivalently, $G$ is complete). It is easy to see that a homogeneous binomial $x_{S_{i_{1}}}\cdots x_{S_{i_{r}}}-x_{S_{j_{1}}}\cdots x_{S_{j_{r}}}\in R[G]$ belongs to $I_{G}$ if and only if $\bigcup_{\ell=1}^{r}S_{i_{\ell}}=\bigcup_{\ell=1}^{r}S_{j_{\ell}}$ as multisets. See, e.g., [3] for details on toric rings and toric ideals.
@@
where $S_{i_{\ell}}=\{j\in[d]\mid G^{(j)}\cap f^{-1}(\ell)\neq\emptyset\}$ for $\ell=1,2,\ldots,k$. Conversely, let $m=x_{S_{i_{1}}}\cdots x_{S_{i_{k}}}\in R[G]$ be a monomial of degree $k$. Then, for ${\mathbf{a}}=(a_{1},\ldots,a_{n})$ with $a_{p}=|\{\ell:p\in S_{i_{\ell}}\}|$, there exists a $k$-coloring $f$ of $G_{{\mathbf{a}}}$ such that ${\mathbf{x}}_{f}=m$ (see [7, Lemma 3.2]). Note that, for $k$-colorings $f$ and $g$ of an induced subgraph of $G$, ${\mathbf{x}}_{f}={\mathbf{x}}_{g}$ if and only if $g$ is obtained from $f$ by permuting colors. It is easy to see that $f\sim_{k}g$ if $g$ is obtained from $f$ by permuting colors. In this paper, we identify $f$ and $g$ if $g$ is obtained from $f$ by permuting colors. Then we can describe a system of generators of $I_{G}$ as follows:
@@
${\mathbf{x}}_{f}-{\mathbf{x}}_{g}\in I_{G}\iff\mbox{$f$ and $g$ are $k$-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and $k\geq\chi(G_{{\mathbf{a}}})$}$
@@
$I_{G}$ $=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are $k$-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and $k\geq\chi(G_{{\mathbf{a}}})$}\rangle$
$=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are k-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in\{0,1,\dots,k\}^{d}$ and $k\geq\chi(G_{{\mathbf{a}}})$}\rangle.$
@@
$\langle[I_{G}]_{2}\rangle$ $=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are 2-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$}\rangle$
$=\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are 2-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in\{0,1,2\}^{d}$}\rangle\subset R[G].$
@@ 3. Gröbner bases
In this section, we give a brief introduction on Gröbner bases. Let $R={\mathbb{K}}[x_{1},x_{2},\ldots,x_{n}]$ be a polynomial ring over a field ${\mathbb{K}}$ with $\deg(x_{i})=1$, and denote $\mathcal{M}_{n}$ the set of all monomials in the variables $x_{1},\ldots,x_{n}$. A monomial order on $R$ is a total order $<$ on $\mathcal{M}_{n}$ such that
@@
Let $u=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$ and $v=x_{1}^{b_{1}}\cdots x_{n}^{b_{n}}$ be monomials in $\mathcal{M}_{n}$. We define the total order $<_{\rm rev}$ on $\mathcal{M}_{n}$ by setting $u<_{\rm rev}v$ if either (i) $\sum_{i=1}^{n}a_{i}<\sum_{i=1}^{n}b_{i}$, or (ii) $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}b_{i}$ and the rightmost nonzero component of the vector $(b_{1}-a_{1},b_{2}-a_{2},\ldots,b_{n}-a_{n})$ is negative. It then follows that $<_{\rm rev}$ is a monomial order on $R$, which is called the (graded) reverse lexicographic order on $R$ induced by the ordering $x_{1}>x_{2}>\cdots>x_{n}$. By reordering the variables, we can obtain another reverse lexicographic order on $R$. Hence, there are $n!$ reverse lexicographic orders on $R$.
Fix a monomial order $<$ on $R$. For a nonzero polynomial $f$ of $R$, the support of $f$, denoted by ${\rm supp}(f)$, is the set of all monomials appearing in $f$ and the initial monomial ${\rm in}_{<}(f)$ of $f$ with respect to $<$ is the largest monomial belonging to ${\rm supp}(f)$ with respect to $<$. Let $I$ be a nonzero ideal of $R$. Then the initial ideal ${\rm in}_{<}(I)$ of $I$ with respect to $<$ is defined as follows:
@@
In general, even if $I=\langle f_{1},\ldots,f_{s}\rangle$, it is not necessarily true that ${\rm in}_{<}(I)=\langle{\rm in}_{<}(f_{1}),\ldots,{\rm in}_{<}(f_{s})\rangle$. A finite set $\{g_{1},\ldots,g_{s}\}$ of nonzero polynomials belonging to $I$ is called a Gröbner basis of $I$ with respect to $<$ if ${\rm in}_{<}(I)=\langle{\rm in}_{<}(g_{1}),\ldots,{\rm in}_{<}(g_{s})\rangle$. Note that if $\{g_{1},\ldots,g_{s}\}$ is a Gröbner basis of $I$, then $I=\langle g_{1},\ldots,g_{s}\rangle$. It is known that $I$ always has a Gröbner basis. A Gröbner basis ${\mathcal{G}}$ for $I$ with respect to $<$ is called reduced if the following conditions hold:
@@
A nonzero ideal $I$ has a unique reduced Gröbner basis with respect to a fixed monomial order $<$. In particular, given a Gröbner basis, we can easily get the reduced Gröbner basis from it.
Next we introduce a method to determine if a finite system of generators of $I$ is a Gröbner basis of $I$ with respect to $<$. For two nonzero polynomials $f$ and $g$ in $R$, the polynomial
@@
Lemma 3.2 (Buchberger Criterion [2, Chapter 2, §9, Theorem 3]).
Let $I$ be a nonzero ideal of $R$, $\mathcal{G}=\{g_{1},\ldots,g_{s}\}$ a finite system of generators of $I$. Then ${\mathcal{G}}$ is a Gröbner basis of $I$ with respect to a monomial order $<$ on $R$ if and only if the remainder of $S$-polynomial $S(g_{i},g_{j})$ on division by ${\mathcal{G}}$ is 00 for all $i\neq j$.
The algorithm called Buchberger Algorithm ([2, Chapter 2, §7, Theorem 2]) is based on Buchberger Criterion, and computes a Gröbner basis for $I$ from a finite system of generators of $I$.
As applications of Gröbner bases, we can determine if a polynomial $f$ belongs to an ideal. This problem is called the ideal membership problem. If ${\mathcal{G}}$ is a Gröbner basis for an ideal $I$ of $R$ with respect to a monomial order $<$, then every polynomial of $f\in R$ has a unique remainder on division by ${\mathcal{G}}$. The remainder is called the normal form of $f$ with respect to ${\mathcal{G}}$.
@@
Let $I$ be a nonzero ideal of $R$ and $\mathcal{G}$ a Gröbner basis of $I$ with respect to a monomial order $<$ on $R$. Fix a polynomial $f$ in $R$. Then $f\in I$ if and only if the normal form of $f$ with respect to ${\mathcal{G}}$ equals 00.
Next, we review a method to compute Hilbert functions. Let $I$ be a graded ideal of $R$. The numerical function $H(R/I,-):{\mathbb{Z}}_{\geq 0}\to{\mathbb{Z}}_{\geq 0}$ with $H(R/I,k)=\dim_{{\mathbb{K}}}R_{k}/I_{k}$ is called the Hilbert function of $R/I$, where $R_{k}$ (resp. $I_{k}$) is the homogeneous part of degree $k$ of $R$ (resp. $I$) and $\dim_{{\mathbb{K}}}R_{k}/I_{k}$ is the dimension of $R_{k}/I_{k}$ as ${\mathbb{K}}$-vector spaces. For a monomial order $<$ on $R$, a monomial $u\in\mathcal{M}_{n}$ is said to be standard with respect to ${\rm in}_{<}(I)$ if $u\notin{\rm in}_{<}(I)$.
@@
Let $I$ be a nonzero graded ideal of $R$ and fix a monomial order $<$ on $R$. Let $\mathcal{B}_{k}$ denote the set of standard monomials of degree $k$ with respect to ${\rm in}_{<}(I)$. Then $\mathcal{B}_{k}$ is a ${\mathbb{K}}$-basis of $S_{k}/I_{k}$ as a ${\mathbb{K}}$-vector space. In particular, one has
@@
$J_{G}:=$ $\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are$2$-colorings of $G_{{\mathbf{a}}}$ with ${\mathbf{a}}\in\{0,1\}^{d}$}\rangle$
$=$ $\langle{\mathbf{x}}_{f}-{\mathbf{x}}_{g}\mid\mbox{$f$ and $g$ are$2$-colorings of an induced subgraph of $G$}\rangle,$
@@
$u_{1}\dots u_{s}-v_{1}\dots v_{s}$
$=$ $u_{2}\cdots u_{s}(u_{1}-v_{1})+u_{3}\cdots u_{s}v_{1}(u_{2}-v_{2})+u_{4}\cdots u_{s}v_{1}v_{2}(u_{3}-v_{3})+\cdots+v_{1}\cdots v_{s-1}(u_{s}-v_{s})$
@@
belongs to $L_{G}$. Since $F$ belongs to $I_{G}$, we have
$\sum_{i=1}^{s}(p_{i}-1)=-s+\sum_{i=1}^{s}p_{i}=-s+\sum_{i=1}^{s}q_{i}=\sum_{i=1}^{s}(q_{i}-1),$
and
@@
Now, we return to the inequalities (4.1). If $G$ is a complete multipartite graph, then $I_{G}$ is generated by quadratic binomials ([6, Theorem 3.1]). Hence from Proposition 4.3 we know that $L_{G}=J_{G}$ implies $\langle[I_{G}]_{2}\rangle=I_{G}$. Therefore, there are the following $6$ cases:

### v2

@@ Abstract.
Kempe equivalence is a classical and important notion on vertex coloring in graph theory. In the present paper, we introduce several ideals associated with graphs and provide a method for determining whether two $k$-colorings are Kempe equivalent via commutative algebra. Moreover, we give a way to compute all $k$-colorings of a graph up to Kempe equivalence by virtue of the algebraic technique on Gröbner bases. As a consequence, the number of $k$-Kempe classes can be computed by using Hilbert functions. Finally, we introduce several algebraic algorithms related to Kempe equivalence.
@@ 1. Introduction
A $k$-coloring $f$ of a graph $G$ on the vertex set $d]:=\{1,2,\ldots,d\}$ is a map from $[d]$ to $[k]$ such that $f(i)\neq f(j)$ for all $\{i,j\}\in E(G)$. The smallest integer $\chi(G)$ such that $G$ has a $\chi(G)$-coloring is called the chromatic number of $G$. Given a $k$-coloring $f$ of $G$, and integers $1\leq i<j\leq k$, let $H$ be a connected component of the induced subgraph of $G$ on the vertex set $f^{-1}(i)\cup f^{-1}(j)$. Then we can obtain a new $k$-coloring $g$ of $G$ by setting
[…]
We say that $g$ is obtained from $f$ by a Kempe switching. Two $k$-colorings $f$ and $g$ of $G$ are called Kempe equivalent, denoted by $f\sim_{k}g$, if there exists a sequence $f_{0},f_{1},\ldots,f_{s}$ of $k$-colorings of $G$ such that $f_{0}=f$, $f_{s}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching. Let denote $\mathcal{C}_{k}(G)$ the set of all $k$-colorings of $G$. Then $\sim_{k}$ is an equivalence relation on $\mathcal{C}_{k}(G)$. The equivalence classes of $\mathcal{C}_{k}(G)$ by $\sim_{k}$ are called the $k$-Kempe classes. We denote ${\rm kc}(G,k)$ the quotient set $\mathcal{C}_{k}(G)/\sim_{k}$ and denote ${\rm Kc}(G,k)$ the number of $k$-Kempe classes of $G$, namely ${\rm Kc}(G,k)=|{\rm kc}(G,k)|$. Kempe switchings were introduced by Kempe in the false proof of the $4$-Color Theorem. However, his idea is powerful in graph coloring theory. Recently, many researchers have studied Kempe switchings and Kempe equivalence. See [5] for an overview of Kempe equivalence.
@@
Let $G$ be a graph on $[d]$ and << a monomial order on $R[G]$, and let $\{{\mathbf{x}}_{f_{1}},\dots,{\mathbf{x}}_{f_{s}}\}$ be the set of all standard monomials of degree $k$ with respect to the initial ideal ${\rm in}_{<}(K_{G})$. Then
@@
The present paper is organized as follows: In Section 2, we will recall the definition of stable set ideals and explain a relationship between stable set ideals and Kempe equivalence. Section 3 will give a brief introduction to Gröbner bases. In Section 4, we will define $2$-coloring ideals and see their algebraic properties. In Section 5, a proof of Theorem 1.1 will be given. In Section 6, we will define Kempe ideals and prove Theorem 1.2. Finally, in Section 7, we will introduce several algorithms related to Kempe equivalence.
@@ 2. Stable set ideals
We regard ${\mathbb{K}}[G]$ as a homogeneous algebra by setting each $\deg({\mathbf{t}}^{\rho(S_{i})}s)=1$. Note that ${\mathbb{K}}[G]$ is a toric ring. Let $R[G]={\mathbb{K}}[x_{S_{1}},\ldots,x_{S_{n}}]$ denote the polynomial ring in $n$ variables over $\mathbb{K}$} with each $\deg(x_{S_{i}})=1$. The stable set ideal of $G$ is the kernel of the surjective homomorphism $\pi:R[G]\to{\mathbb{K}}[G]$ defined by $\pi(x_{S_{i}})={\mathbf{t}}^{\rho(S_{i})}s$ for $1\leq i\leq n$. Note that $I_{G}$ is a toric ideal, and hence a prime ideal generated by homogeneous binomials. The toric ring ${\mathbb{K}}[G]$ is called quadratic if $I_{G}$ is generated by quadratic binomials. We say that $I_{G}$ is generated by quadratic binomials” even if $I_{G}=\{0\}$ (or equivalently, $G$ is complete). It is easy to see that a homogeneous binomial $x_{S_{i_{1}}}\cdots x_{S_{i_{r}}}-x_{S_{j_{1}}}\cdots x_{S_{j_{r}}}\in R[G]$ belongs to $I_{G}$ if and only if $\bigcup_{\ell=1}^{r}S_{i_{\ell}}=\bigcup_{\ell=1}^{r}S_{j_{\ell}}$ as multisets. See, e.g., [3] for details on toric rings and toric ideals.
@@
where $S_{i_{\ell}}=\{j\in[d]\mid G^{(j)}\cap f^{-1}(\ell)\neq\emptyset\}$ for $\ell=1,2,\ldots,k$. Conversely, let $m=x_{S_{i_{1}}}\cdots x_{S_{i_{k}}}\in R[G]$ be a monomial of degree $k$. Then, for ${\mathbf{a}}=(a_{1},\ldots,a_{n})$ with $a_{p}=|\{\ell:p\in S_{i_{\ell}}\}|$, there exists a $k$-coloring $f$ of $G_{{\mathbf{a}}}$ such that ${\mathbf{x}}_{f}=m$ (see [7, Lemma 3.2]). For example, we consider the graphs $G$ and $G_{{\mathbf{a}}}$ with ${\mathbf{a}}=(2,1,0,2)$ as follows:
11224433
(a) $G$
$1_{1}$$1_{2}$$2_{1}$$4_{1}$$4_{2}$
(b) $G_{{\mathbf{a}}}$ with ${\mathbf{a}}=(2,1,0,2)$
We define a $4$-coloring $f$ of $G_{{\mathbf{a}}}$ by
$f(i)=\begin{cases}1&i\in\{1_{1},4_{1}\},\\ 2&i\in\{2_{1}\},\\ 3&i\in\{1_{2}\},\\ 4&i\in\{4_{2}\}.\end{cases}$
Then since $S_{i_{1}}=\{1,4\},S_{i_{2}}=\{2\},S_{i_{3}}=\{1\},S_{i_{4}}=\{4\}$, one has
${\mathbf{x}}_{f}=x_{\{1,4\}}x_{\{2\}}x_{\{1\}}x_{\{4\}}.$
On the other hand, we can obtain the $4$-coloring $f$ (up to exchanging colors and exchanging the coloring of vertices in each clique $G^{(i)}$ of $G_{\mathbf{a}}$) from the monomial $x_{\{1,4\}}x_{\{2\}}x_{\{1\}}x_{\{4\}}$ as follows: for each variable $x_{S}$, we assign one color to a single copy $i_{j}$ of each vertex $i\in S$.
Note that, for $k$-colorings $f$ and $g$ of an induced subgraph of $G$, ${\mathbf{x}}_{f}={\mathbf{x}}_{g}$ if and only if $g$ is obtained from $f$ by permuting colors. It is easy to see that $f\sim_{k}g$ if $g$ is obtained from $f$ by permuting colors. In this paper, we identify $f$ and $g$ if $g$ is obtained from $f$ by permuting colors. Then we can describe a system of generators of $I_{G}$ as follows:
@@
[…]
@@
[…]
@@
[…]
@@ 3. Gröbner bases
In this section, we give a brief introduction to Gröbner bases. Let $R={\mathbb{K}}[x_{1},x_{2},\ldots,x_{n}]$ be a polynomial ring over a field $\mathbb{K}$} with $\deg(x_{i})=1$, and denote $\mathcal{M}_{n}$ the set of all monomials in the variables $x_{1},\ldots,x_{n}$. A monomial order on $R$ is a total order << on $\mathcal{M}_{n}$ such that
@@
Let $u=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$ and $v=x_{1}^{b_{1}}\cdots x_{n}^{b_{n}}$ be monomials in $\mathcal{M}_{n}$. We define the total order $_{\rm rev}$ on $\mathcal{M}_{n}$ by setting $u<_{\rm rev}v$ if either (i) $\sum_{i=1}^{n}a_{i}<\sum_{i=1}^{n}b_{i}$, or (ii) $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}b_{i}$ and the rightmost nonzero component of the vector $(b_{1}-a_{1},b_{2}-a_{2},\ldots,b_{n}-a_{n})$ is negative. It then follows that $_{\rm rev}$ is a monomial order on $R$, which is called the (graded) reverse lexicographic order on $R$ induced by the ordering $x_{1}>x_{2}>\cdots>x_{n}$. By reordering the variables, we can obtain another reverse lexicographic order on $R$. Hence, there are $n!$ reverse lexicographic orders on $R$.
Fix a monomial order << on $R$. For a nonzero polynomial $f$ of $R$, the support of $f$, denoted by ${\rm supp}(f)$, is the set of all monomials appearing in $f$ and the initial monomial ${\rm in}_{<}(f)$ of $f$ with respect to << is the largest monomial belonging to ${\rm supp}(f)$ with respect to <<. Let $I$ be a nonzero ideal of $R$. Then the initial ideal ${\rm in}_{<}(I)$ of $I$ with respect to << is defined as follows:
@@
In general, even if $I=\langle f_{1},\ldots,f_{s}\rangle$, it is not necessarily true that
${\rm in}_{<}(I)=\langle{\rm in}_{<}(f_{1}),\ldots,{\rm in}_{<}(f_{s})\rangle.$
A finite set $\{g_{1},\ldots,g_{s}\}$ of nonzero polynomials belonging to $I$ is called a Gröbner basis of $I$ with respect to << if ${\rm in}_{<}(I)=\langle{\rm in}_{<}(g_{1}),\ldots,{\rm in}_{<}(g_{s})\rangle$. Note that if $\{g_{1},\ldots,g_{s}\}$ is a Gröbner basis of $I$, then $I=\langle g_{1},\ldots,g_{s}\rangle$. It is known that $I$ always has a Gröbner basis. A Gröbner basis $\mathcal{G}$} for $I$ with respect to << is called reduced if the following conditions hold:
@@
A nonzero ideal $I$ has a unique reduced Gröbner basis with respect to a fixed monomial order <<. In particular, given a Gröbner basis, we can easily get the reduced Gröbner basis from it.
Next we introduce a method for determining whether a finite system of generators of $I$ is a Gröbner basis of $I$ with respect to <<. For two nonzero polynomials $f$ and $g$ in $R$, the polynomial
@@
Lemma 3.2 (Buchberger’s Criterion [2, Chapter 2, §9, Theorem 3]).
Let $I$ be a nonzero ideal of $R$, $\mathcal{G}=\{g_{1},\ldots,g_{s}\}$ a finite system of generators of $I$. Then $\mathcal{G}$} is a Gröbner basis of $I$ with respect to a monomial order << on $R$ if and only if the remainder of $S$-polynomial $S(g_{i},g_{j})$ on division by $\mathcal{G}$} is 0 for all $i\neq j$.
The algorithm called Buchberger’s Algorithm ([2, Chapter 2, §7, Theorem 2]) is based on Buchberger’s Criterion, and computes a Gröbner basis for $I$ from a finite system of generators of $I$.
As applications of Gröbner bases, we can determine if a polynomial $f$ belongs to an ideal. This problem is called the ideal membership problem. If $\mathcal{G}$} is a Gröbner basis for an ideal $I$ of $R$ with respect to a monomial order <<, then every polynomial of $f\in R$ has a unique remainder on division by $\mathcal{G}$}. The remainder is called the normal form of $f$ with respect to $\mathcal{G}$}.
@@
Let $I$ be a nonzero ideal of $R$ and $\mathcal{G}$ a Gröbner basis of $I$ with respect to a monomial order << on $R$. Fix a polynomial $f$ in $R$. Then $f\in I$ if and only if the normal form of $f$ with respect to $\mathcal{G}$} equals 0.
Next, we review a method for computing Hilbert functions. Let $I$ be a graded ideal of $R$. The numerical function $H(R/I,-):{\mathbb{Z}}_{\geq 0}\to{\mathbb{Z}}_{\geq 0}$ with $H(R/I,k)=\dim_{{\mathbb{K}}}R_{k}/I_{k}$ is called the Hilbert function of $R/I$, where $R_{k}$ (resp. $I_{k}$) is the homogeneous part of degree $k$ of $R$ (resp. $I$) and $\dim_{{\mathbb{K}}}R_{k}/I_{k}$ is the dimension of $R_{k}/I_{k}$ as $\mathbb{K}$}-vector spaces. For a monomial order << on $R$, a monomial $u\in\mathcal{M}_{n}$ is said to be standard with respect to ${\rm in}_{<}(I)$ if $u\notin{\rm in}_{<}(I)$.
@@
Let $I$ be a nonzero graded ideal of $R$ and fix a monomial order << on $R$. Let $\mathcal{B}_{k}$ denote the set of standard monomials of degree $k$ with respect to ${\rm in}_{<}(I)$. Then $\mathcal{B}_{k}$ is a $\mathbb{K}$}-basis of $R_{k}/I_{k}$ as a $\mathbb{K}$}-vector space. In particular, one has
@@
[…]
@@
$u_{1}\dots u_{s}-v_{1}\dots v_{s}$ $=$ $u_{2}\cdots u_{s}(u_{1}-v_{1})+u_{3}\cdots u_{s}v_{1}(u_{2}-v_{2})+u_{4}\cdots u_{s}v_{1}v_{2}(u_{3}-v_{3})$
$+\cdots+v_{1}\cdots v_{s-1}(u_{s}-v_{s})$
@@
belongs to $L_{G}$. Since $F$ belongs to $I_{G}$, $\pi(\prod_{i=1}^{s}x_{S_{i}})=\pi(\prod_{i=1}^{s}x_{S_{i}^{\prime}})$. This implies that $\bigcup_{i=1}^{s}S_{i}=\bigcup_{i=1}^{s}S^{\prime}_{i}$ as multisets. Hence one has $\sum_{i=1}^{s}p_{i}=\sum_{i=1}^{s}q_{i}$ and
@@
Now, we return to the inclusions (4.1). If $G$ is a complete multipartite graph, then $I_{G}$ is generated by quadratic binomials ([6, Theorem 3.1]). Hence from Proposition 4.3 we know that $L_{G}=J_{G}$ implies $\langle[I_{G}]_{2}\rangle=I_{G}$. Therefore, there are the following $6$ cases: